R=(10)log(I/Io), where R=90. Compare I to Io.
Answer by stanbon(57323)

R=(10)log(I/Io), where R=90. Compare I to Io.
-------
Solve: 10*log(I/Io) = 90
-----
log(I/Io) = 9
---
I/Io = 10^9
---
I/Io = 1,000,000,000
=======================
Cheers,
Stan H.